# BoostTrack++: using tracklet information to detect more objects in multiple object tracking

### v1

@@ 5.1 Datasets and metrics
Datasets. We perform experiments and evaluation of our method on standard MOT benchmark datasets: MOT17  [16] and MOT20  [17]. MOT17 contains pedestrian videos filmed with static and moving camera. The training set consists of $7$ sequences and $5316$ frames in total (frame rate ranges from $14$ to $30$ depending on the video), while test set contains continuation of the same sequences and has $5919$ frames in total.
@@ 5.3 Ablation study
Since the DLO confidence boost resulted in substantial performance improvement on the MOT20, but also in an increased number of IDs and IDWSs, we study the effect of various similarity measures from equation  10 on MOT20 validation set. We performed experiments on three baselines: SORT+DLO, BoostTrack and BoostTrack+. Table  2 shows the results [^5] We consider the best IDs and IDWS values that are the closest to non DLO baseline. of using SORT+DLO and BoostTrack+ as baselines (as the most basic and the more advanced methods, respectively). We display results of using the BoostTrack baseline in  A.
The first two rows of table  2 are the baselines. The first row shows results of the baseline method without DLO confidence boost, while the second uses simple IoU based DLO confidence boost which we try to improve (see equation  4). A successful DLO method should keep IDs and IDSW values as close to the values from the first row, and at the same time improve the HOTA, MOTA and IDF1 values from the second row.
Table 2: The effect of various similarity measures used for DLO confidence boost on MOT20 validation set (best in bold).
@@ 5.3.2 Influence of other components
We study the influence of proposed components on BoostTrack+ baseline on MOT17 and MOT20 validation sets. In  B we show results of the ablation study where we used SORT  [31] and BoostTrack as baselines.
Table  3 shows the influence of adding each of the proposed component to the BoostTrack+ baseline: namely, new similarity measure for the confidence boost (S column), soft (detection confidence) boost (SB column), and using varying threshold (VT column) for the DLO confidence boost. Note, when we use SB or VT, and not S, the similarity used in SB and VT is IoU. For more details, see algorithm  1. Again, in the first row, we show results of the baseline method without any DLO confidence boost, while the second row shows results of the original DLO confidence boost from BoostTrack.
@@
The results from table  3 show that using SB and VT with IoU as a similarity measure (without the proposed average similarity) increases the number of IDs even more than baseline DLO. However, combined with S, we get the best of both worlds: we get a substantial improvement in tracking performance, outperforming the original DLO confidence boost, while at the same time only slightly increasing the number of IDs and IDSWs: +5 IDSWs (6.3%), +8 IDs (2%) on the MOT17 and +2 IDSWs (0.3% ), +54 IDs (2.9%) on the MOT20.
@@ 5.4 Comparison with other methods
To evaluate our method on MOT17  [16] and MOT20  [17] test sets (under private detection protocol), we used S+SB+VT setting. The results on the test sets are provided in table  4. We mark offline methods with ’*’666 Note that our method, as most ”online” methods (e.g.  [9, 10, 13, 11, 35]), applies postprocessing to the results (which are obtained online), and could be best described as ”semi-online”. However, the official MOT Challenge  [43] distinguishes only offline and online methods, and methods like ours are considered online. .
@@
Our BoostTrack++ method achieves improvement compared to the BoostTrack+ on MOT17 test set: +0.2 HOTA, +0.1 MOTA, +0.4 IDF1, -24 IDSWs. On more challenging MOT20, BoostTrack++ shows slightly better improvement compared to the BoostTrack+: +0.2 HOTA, +0.5 MOTA, +0.5 IDF1, -65 IDSWs. Improvement in MOTA indicates the increase in the use of true positive detections, which improves HOTA and IDF1 scores.
@@ 6 Conclusions
Appendix A Influence of different similarity measures
In table  5 we show results of applying various similarity measures instead of IoU for the DLO confidence boost. We used BoostTrack as the baseline method. As with SORT or BoostTrack+ baselines (displayed in table  2), using the average of all three similarity measures provides the best trade-off between introducing additional identities and overall tracking performance.
@@ Appendix A Influence of different similarity measures
Appendix B Ablation study on SORT and BoostTrack
We also performed an ablation study using SORT and BoostTrack+ as baseline methods. Table  6 shows results of various settings on MOT20 validation set. Using only VT with BoostTrack+ as baseline provides a good example of trade-offs required. It even reduced IDs and IDWSs by 2. However, the overall tracking performance is only slightly improved, while S or S+SB+VT settings provide a better trade-off between various metrics and the number of additional IDs and IDSWs.
@@ Appendix B Ablation study on SORT and BoostTrack
Setting SORT baseline BoostTrack+ baseline

### v2

@@ 5.1 Datasets and metrics
Datasets. We perform experiments and evaluation of our method on standard MOT benchmark datasets: MOT17  [16] and MOT20  [17]. MOT17 contains pedestrian videos filmed with static and moving camera. The training set consists of $7$ sequences and $5316$ frames in total (frame rate ranges from $14$ to $30$ depending on the video), while the test set contains a continuation of the same sequences and has $5919$ frames in total.
@@ 5.3 Ablation study
Since the DLO confidence boost resulted in substantial performance improvement on the MOT20, but also in an increased number of IDs and IDWSs, we study the effect of various similarity measures from equation  10 on MOT20 validation set. We performed experiments on three baselines: SORT+DLO, BoostTrack and BoostTrack+. Table  3 shows the results [^6] We consider the best IDs and IDWS values that are the closest to non-DLO baseline. of using SORT+DLO and BoostTrack+ as baselines (as the most basic and the more advanced methods, respectively). We display the results of using the BoostTrack baseline in  B.
The first two rows of table  3 are the baselines. The first row shows results of the baseline method without DLO confidence boost, while the second uses a simple IoU based DLO confidence boost which we try to improve (see equation  4). A successful DLO method should keep IDs and IDSW values as close to the values from the first row, and at the same time improve the HOTA, MOTA and IDF1 values from the second row.
Table 3: The effect of various similarity measures used for DLO confidence boost on the MOT20 validation set (best in bold).
@@ 5.3.2 Influence of other components
We study the influence of proposed components on BoostTrack+ baseline on MOT17 and MOT20 validation sets. In  C we show the results of the ablation study where we used SORT  [31] and BoostTrack as baselines.
Table  4 shows the influence of adding each of the proposed components to the BoostTrack+ baseline: namely, new similarity measure for the confidence boost (S column), soft (detection confidence) boost (SB column), and using varying threshold (VT column) for the DLO confidence boost. Note, when we use SB or VT, and not S, the similarity used in SB and VT is IoU. For more details, see algorithm  1. Again, in the first row, we show results of the baseline method without any DLO confidence boost, while the second row shows results of the original DLO confidence boost from BoostTrack.
@@
The results from table  4 show that using SB and VT with IoU as a similarity measure (without the proposed average similarity) increases the number of IDs even more than baseline DLO. However, combined with S, we get the best of both worlds: we get a substantial improvement in tracking performance, outperforming the original DLO confidence boost, while at the same time only slightly increasing the number of IDs and IDSWs: +5 IDSWs (6.3%), +8 IDs (2%) on the MOT17 and +2 IDSWs (0.3% ), +54 IDs (2.9%) on the MOT20.
@@ 5.4 Comparison with other methods
To evaluate our method on MOT17  [16] and MOT20  [17] test sets (under private detection protocol), we used all proposed additions combined (S+SB+VT setting). The results on the test sets are provided in the table  5. We mark offline methods with ’*’777 Note that our method, as most ”online” methods (e.g.  [9, 10, 13, 11, 35]), applies postprocessing to the results (which are obtained online), and could be best described as ”semi-online”. However, the official MOT Challenge  [43] distinguishes only offline and online methods, and methods like ours are considered online. .
@@
Our BoostTrack++ method achieves improvement compared to the BoostTrack+ on the MOT17 test set: +0.2 HOTA, +0.1 MOTA, +0.4 IDF1, -24 IDSWs. On more challenging MOT20, BoostTrack++ shows slightly better improvement compared to the BoostTrack+: +0.2 HOTA, +0.5 MOTA, +0.5 IDF1, -65 IDSWs. Improvement in MOTA indicates the increase in the use of true positive detections, which improves HOTA and IDF1 scores.
@@ 6 Conclusions
Appendix A List of abbreviations
In table  6, we show the list of all abbreviations used in the paper.
Table 6: List of abbreviations
Abbreviation Definition
TBD MOT Tracking by detection multiple object tracking
IoU Intersection over union
BIoU Buffered IoU from  [15]
SBIoU Soft BIoU introduced in subsection  4.1
DLO Detection of likely objects used to boost confidence scores in  [14]
IDSW Identity switch
S (ablation study) Similarity measure from equation  10
SB (ablation study) Soft (detection confidence) boost from subsection  4.3
VT (ablation study) Varying similarity threshold from subsection  4.4
Appendix B Influence of different similarity measures
In table  7 we show the results of applying various similarity measures instead of IoU for the DLO confidence boost. We used BoostTrack as the baseline method. As with SORT or BoostTrack+ baselines (displayed in table  3), using the average of all three similarity measures provides the best trade-off between introducing additional identities and overall tracking performance.
@@ Appendix B Influence of different similarity measures
Appendix C Ablation study on SORT and BoostTrack
To show the robustness of the proposed components, we performed additional ablation experiments using SORT and BoostTrack as baseline methods. Table  8 shows the results of various settings on MOT20 validation set. Note that, ideally, our method should not introduce new IDs and IDSWs, and it should at the same time improve overall tracking performance. Using only VT with BoostTrack as the baseline provides a good example of trade-offs required. Adding VT even reduced IDs and IDWSs by 2. However, the overall tracking performance is only slightly improved by adding VT only, while S or S+SB+VT settings provide a better trade-off between various metrics and the number of additional IDs and IDSWs.
@@ Appendix C Ablation study on SORT and BoostTrack
Setting SORT baseline BoostTrack baseline